you know that about 25% of customers will call back within 48 hours. of course customers who call back within 48 hours, about 20% of them will call back within 72 hours. of the 120 people you spoke with today how many do you expect to call two more times in the next 5 days
1
Expert's answer
2020-05-12T16:08:42-0400

To solve this problem, we will start with "120" people.

Since we know that about "25 \%" of people will call back within "48" hours.

So:

"25 \% \, \text{of} \, 120 \\
= \frac{25}{100} \times 120 \\
= 30"

Hence "30" people out of "120" people will call back within "48" hours

Again:

Of those "30" people, we know that about "20 \%" of them will call back within "72" hours.

So:

"20 \% \, \text{of} \, 30 \\
= \frac{20}{100} \times 30 \\
= 6"

Hence:

"6" people will call two more times in the next "5" days.